Solve: 3 ( x- 1/3) = 7(x-1/7)
Answer by Alan3354(69443)

Solve: 3 ( x- 1/3) = 7(x-1/7)
----------------
get rid of parentheses:
3x - 1 = 7x - 1
3x = 7x
x = 0 (or infinity)